Al, Ronda, and Rony enter into a partnership. Al owns 1/4 of the company and Ronda 1/8. Calculate what part is owned by Rony.
Answer by SHUgrad05(58)

Al= 1/4 or 4/8
Ronda= 1/8
Rony= ? (lets call this x)
4/8 + 1/8 + x = 8/8
5/8 + x= 8/8
x= 3/8
Rony owns 3/8 of the company